Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2014-03-01

Observers: Heather Campbell, Nadejda Blagorodnova
Support team: Morgan Fraser
TAT: Joe Anderson, Maria Teresa Botticella

Weather:

Moon down during all night
Beginning night: Humidity very high started about 60%, and rose to almost 80%, wind ~3m/s, DIMM seeing 1", no signs of clouds.

Middle of the night: Wind increased a little ~5-10m/s. Humidity dropped to ~50% at about 4UT. DIMM seeing~1-1.2'.

End of the night: Wind ~5-10m/s. Humidity rising again to ~70%. DIMM seeing ~1'.

Observations:

Started the night with SOFI.

0.00 Image analysis on SOFI. Failed - due to the high humidity.

        TO tried just a focus. But no better as humidity rising.

        TO went to telescope to measure humidly. 

0.50 Closed Dome - condensation on the mirror. Turned on the lamp to dry the mirror.

1.10 Humidity 75% and still rising.

1.50 Opened dome again. And Image analysis, still not good.

2.15 Started 2013ddu followup. But images looked very bad, seeing about 3'. 

2.25 Aborted and redid image analysis to check the mirror etc, still looked bad.

2.35 Gave up using SOFI, humidity too high for useful SOFI imaging and switched to EFOSC. Unfortunately now these followup observations cannot be done during this subrun.

Changed to EFOSC

2.50 Focus and image analysis.

3.05 SKY_J5271481-4716073 1500s exposure. Low signal to noise, clear emission line at z=0.09.

3.40  LTT745a spectrophotometric standard.

4.05 MLS140220-081342+181306  1200s exposure. Very low S/N, slightly blue continuum. Dropped to low to return to tonight. Still visible next subrun.

4.35 LSQ14abd 900s exposure. Normal Ia, z~0.097, prepeak.

5.00 OGLE-2014-SN-024 1500s exposure. SNe Ia (best fits to Ia-91T), z~0.1, prepeak.

5.35 LSQ14abu 900s exposure. Fairly low S/N, not well fit by snid. A black body temperature gives ~5300K, which is similar to a K star and the spectra resembles it (He spectral features, including absorption in H gamma and 5182.604 Mg1). Seams to have moved ~0.15arc min from the LSQ position on 28/2/2014. Could this be a something interesting?

6.05 SN2013bb. Took 100s acquisition image in R to see target. Followup 2*1200s to complete the rest of the followup started last night.

7.35 PS1-14nt. 1500s exposure. Type II ~30-40 days post peak z~0.04. Seeing improving to ~1'.

8.05 LSQ14aab. 900s exposure. Very Low signal to noise. Blueish continuum.

8.35 PS1-ny. 900s exposure. Normal SNe Ia, z~0.88, 1-2weeks post peak.

8.55 LSQ14zl. 1000s exposure.  Looked like a CV, H gamma and  H delta absorption lines.

Twilight 9.12

9.20 Standard EG274

9.45 LSQ14uv. 500s exposure. Looks like an M star.

10.00 Skyflats gr  (z dropped as not used in this subrun and had problems yesterday trying to take it)






Comments